can you help me solve: f(x)= x-4 and g(x)=x+2
1. f-g 2. f*g

Answer by stanbon(75887)

can you help me solve: f(x)= x-4 and g(x)=x+2
1. f-g
=(f-g)(x) = f(x) - g(x)
= (x-4) - (x+2)
= -6
-------------------
2. f*g
= f(x)*g(x)
=(x-4)(x+2)
=x^2-2x-8
-----------------
Cheers,
Stan H.